OpenStudy (curry):

find an equivalent equation in rectangular form for r-6sin=2coss

OpenStudy (anonymous):

do you know the substitutions?

OpenStudy (curry):

ye i got to x^2 +y^2=2x+6y but i dont know where to go from there2

OpenStudy (experimentx):

put sin .. = y/sqrt(x^2+y^2) and cos ..= x/sqrt(x^2+y^2) and r = sqrt(x^2+y^2) and then simplify

OpenStudy (curry):

but the final ansswer should be in rectangular form meaning it hould have no r or anythuing

OpenStudy (experimentx):

yeah ... that's why i am telling you to substitute r's and sines and cosines ... as above

OpenStudy (curry):

but either way i still end up with x^2 + y^2 = 2x+ 6y can i plz see ur woprk and final anwer

OpenStudy (anonymous):

your answer is correct. what do you need?

OpenStudy (anonymous):

your answer can be written in standard form for a circle but that's it...

OpenStudy (anonymous):

is that what you want to do?

OpenStudy (curry):

ye

OpenStudy (curry):

y would it equal 3?

OpenStudy (curry):

i mean t

OpenStudy (anonymous):

ok.. let's start with your equation: x^2 + y^2 = 2x + 6y

OpenStudy (anonymous):

move everything on that right side over to the left side. can you do that?

OpenStudy (curry):

ye

OpenStudy (anonymous):

so what do you have?

OpenStudy (curry):

x(x-2)=-y(y-6)

OpenStudy (anonymous):

don't factor...

OpenStudy (anonymous):

and your right side should be 0

OpenStudy (curry):

x^2 - 2x + y^2 -6y

OpenStudy (anonymous):